A grating that has 3400 slits per cm produces a third-order fringe at a 24.0? angle.

What wavelength of light is being used?
Answer #1

the gap between the slits by the formula d=1/N


N is number of lines per metre so 3400 line/cm = 340000 lines/m

So d=2.941*10^-6


Now we use the formula dsin(angle)=n(wavelength)


d is the same

n is the order of the diffraction pattern


so wavelenth=dsin(angle)/n

=[(2.941*10^-6)*sin24]/3

=3.9876*10^-7 m